# Curriculum-Guided Layer Scaling for Language Model Pretraining

### v2

@@ A.1 Experimental Setup
A.2 Few Shot Results
In standard academic pretraining works, few-shot evaluations are uncommon because effective in-context learning emerges at larger scales, as shown in the GPT-3 paper Brown et al. (2020). For example, recent large-scale academic works such as H-Nets Hwang et al. (2025), which pretrains 1B parameter models for 100B tokens, do not include few-shot evaluations. Nevertheless, we conduct some few-shot evaluations, with PMI accuracy Gu et al. (2024) to reduce variance, and report the results in Table 6. For both zero- and 5-shot evaluation, CGLS achieves the best average across the benchmarks.
@@ A.4 Stability
A.5 Perplexity Results

### v3

@@ A.1 Experimental Setup
A.2 Few Shot Results
In standard academic pretraining works, few-shot evaluations are uncommon because effective in-context learning emerges at larger scales, as shown in the GPT-3 paper (Brown et al., 2020). For example, recent large-scale academic works such as H-Nets (Hwang et al., 2025), which pretrains 1B parameter models for 100B tokens, do not include few-shot evaluations. Nevertheless, we conduct some few-shot evaluations, with PMI accuracy (Gu et al., 2024) to reduce variance, and report the results in Table 6. For both zero- and 5-shot evaluation, CGLS achieves the best average across the benchmarks.
A.3 Hyperparameter Ablations
@@ A.4 Stability
A.5 Robustness to the Difficulty Classifier
Model PIQA ARC-E ARC-C HellaSwag Lambada Average
Llama-3.2-1B w/ 2.5B Tokens
Baseline 59.09 36.36 24.24 34.20 32.99 37.38
CGLS 61.21 41.37 25.43 34.65 32.74 39.08
FineWeb 61.59 45.20 26.37 31.30 30.84 39.06
Table 8: Comparison of CGLS using the default DistilBERT-based difficulty classifier vs. a FineWeb-based difficulty classifier. All benchmarks are conducted in a zero-shot setting. FineWeb yields similar average performance to standard CGLS, with stronger results on ARC-E and ARC-C but lower performance on HellaSwag and Lambada.
A core component of CGLS is the curriculum used to order training data by difficulty. In our primary experiments, difficulty stratification is performed using a DistilBERT-based classifier trained to predict coarse complexity labels. Since model performance may depend on the quality or structure of this classifier, we assess the robustness of CGLS to alternate curriculum signals.
To this end, we conducted an additional experiment in which we replaced the DistilBERT-based classifier with the FineWeb educational difficulty classifier(Penedo et al., 2024), which provides a substantially different distributional prior and difficulty spectrum. Table 8 reports results for the baseline, the original CGLS configuration, and the variant that uses FineWeb for curriculum construction.
Across benchmarks, the FineWeb-based CGLS variant performs comparably to the DistilBERT-based version: both substantially outperform the baseline, and their average accuracies (39.06% vs. 39.08%) are nearly identical. The FineWeb classifier produces modest improvements on PIQA, ARC-Easy, and ARC-Challenge, while slightly reducing performance on HellaSwag and Lambada. These shifts reflect differences in how the two classifiers stratify training data, but the overall stability indicates that CGLS is not overly sensitive to the specific classifier used to define difficulty.
These findings suggest that CGLS benefits from the existence of a meaningful progression in data complexity, rather than from the idiosyncrasies of a particular classifier. We therefore view improved, domain-adaptive, or multi-signal difficulty estimators as an important direction for future work, and emphasize that CGLS is compatible with a wide family of curriculum-generation strategies beyond DistilBERT.
A.6 Perplexity Results